18.10(b) If 175 cm of carbon dioxide effuses through a small hole in 255 seconds, how long wil it take the same volume of sulfur dioxide to effuse under the same temperature and pressure?
Answer #1

To calculate the relative rate, of effusion of CO2 and SO2 gas uses following equation Rate of effusion of Co Rate of effusio

r1/r2 = (v1/t1) / (v2/t2) = square root of (M2/M1)
where r1 is rate of effussion of gas CO2

r2 is rate of effussion of gas SO2
v1 = volume of gas effussed for gas CO2
v2 = volume of gas effussed for gas SO2
t1 = time taken by gas CO2
t2 = time taken by gas SO2

Then;

r1/r2 = (v1/t1) / (v2/t2)

1.21 =( 175 cm^3/ 255 seconds ) / (175 cm^3/ t2 seconds )

t2 seconds = 1.21*255

= 308.55 seconds